# Probing Neural Topology of Large Language Models

### v2

@@ A.7 Graph probing on vision language models
Unlike LLMs, which contain only textual hidden states, Vision-Language Models (VLMs) jointly encode image and text features within a shared transformer backbone, producing multimodal hidden states that remain underexplored. To study their internal topology, we extract the hidden representations at each layer and compute correlations as in Equations (1-3), yielding a connectivity matrix whose co-activation time series span both modalities.
We evaluated neural topology in VLMs by probing different sizes of LLaVA-v1.5 (Liu et al., 2024). For this setup, we adapted the CLEVR dataset (Johnson et al., 2016) into an object-counting classification task of 10,000 randomly sampled images with labels corresponding to the number of objects (integers 3-10). GNN-based probes (Equations (14-16)) were trained with cross-entropy loss on neural topology graphs constructed at a fixed density of 0.01. As shown in Figure 9(a), graph probing consistently outperformed linear activation probing across both the 7B and 13B models, validating that neural topology is more informative than neural activation in VLMs regarding visual understanding capabilities.

### v3

@@ A.7 Graph probing on vision language models
Unlike LLMs, which contain only textual hidden states, Vision-Language Models (VLMs) jointly encode image and text features within a shared transformer backbone, producing multimodal hidden states that remain underexplored. To study their internal topology, we extract the hidden representations at each layer and compute correlations as in Equations (1-4), yielding a connectivity matrix whose co-activation time series span both modalities.
We evaluated neural topology in VLMs by probing different sizes of LLaVA-v1.5 (Liu et al., 2024). For this setup, we adapted the CLEVR dataset (Johnson et al., 2016) into an object-counting classification task of 10,000 randomly sampled images with labels corresponding to the number of objects (integers 3-10). GNN-based probes (Equations (12-14)) were trained with cross-entropy loss on neural topology graphs constructed at a fixed density of 0.01. As shown in Figure 9(a), graph probing consistently outperformed linear activation probing across both the 7B and 13B models, validating that neural topology is more informative than neural activation in VLMs regarding visual understanding capabilities.